OpenStudy (anonymous):

What is the limit of the series 1 +2/5 +4/25 + 8/125 ???????????????????????????????????????

OpenStudy (kanwal32):

is it infinite series

OpenStudy (kanwal32):

the sum would be

OpenStudy (kanwal32):

\[\frac{ 1}{ 1-2/5 }\]

OpenStudy (anonymous):

ohhhhhhhhhhhhhhhhhhhhhhhhhhh I can't beleive I forgot that thanks

OpenStudy (anonymous):

wait is this the formula: 

OpenStudy (anonymous):

\[\sum _{n=0}^{\infty } \frac{2^n}{5^n}=\frac{5}{3} \]